実用外国語演習 (英語) 2 単位 (選択) 3 年 (後期)

実用外国語演習 (英語)  座喜純 · 准教授/ 大学院ソシオ・アーツ・アンド・サイエンス研究部

【授業目的】This class formed to give students exposure to and practice with the English writing or speech-making process from collecting ideas to the final version. Students will learn how to build paragraphs and outlines, and proof-reading for English essays or speech presentations.

【授業概要】 English composition and speech presentation 【キーワード】 Listening， Speaking， Writing 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】 By the end of this course, students should be able to write and present

competently on almost any topic. 【授業計画】

1. General Class Guidance

2. Topic 1

3. Presentation

4. Topic 1

5. Presentation

6. Topic 1

7. Presentation

8. Topic 1

9. Presentation

10. Topic 1

11. Presentation

12. Topic 1

13. Presentation

14. Topic 1

15. Presentation 【成績評価】 ) weekly attendance, participation to class work, and homework are

very important 【教科書】 [教科書] 【参考書】 [参考資料] 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220083 【連絡先】

⇒ 座喜 . (オフィスアワー: contact: jzaki@ias.tokushima-u.ac.jp)

実用外国語演習 (英語) 2 units (selection) 3rd-year(2nd semester)

Jun Zaki · Associate Professor / Institute of Socio-Arts and Sciences

Target⟩ This class formed to give students exposure to and practice with the English writing or speech-making process from collecting ideas to the final version. Students will learn how to build paragraphs and outlines, and proof-reading for English essays or speech presentations.

Outline⟩ English composition and speech presentation �������⟩ Listening, Speaking, Writing Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ By the end of this course, students should be able to write and present

competently on almost any topic.

Schedule⟩

1. General Class Guidance

2. Topic 1

3. Presentation

4. Topic 1

5. Presentation

6. Topic 1

7. Presentation

8. Topic 1

9. Presentation

10. Topic 1

11. Presentation

12. Topic 1

13. Presentation

14. Topic 1

15. Presentation Evaluation Criteria⟩ ) weekly attendance, participation to class work, and

homework are very important Textbook⟩ [教科書] Reference⟩ [参考資料] Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220083 Contact⟩

⇒ Zaki . (Office Hour: contact: jzaki@ias.tokushima-u.ac.jp)